# Extreme Compression of Large Language Models via Additive Quantization

### 11 Jan 2024 v1

@@ 3.1 Overview
Our algorithm starts from the intuition that additive quantization (AQ) aims to solve a similar problem to standard approaches in post-training quantization (PTQ) (Nagel et al., 2020; Frantar et al., 2022b): specifically, both settings assume the existence of a set of “input” vectors, i.e. input data for AQ, and the weight matrix rows for PTQ. The goal is to compress these inputs while preserving dot product similarity, against query vectors (for AQ), and against layer input embeddings (for PTQ). The difference between the two is that, AQ normally assumes that the distribution of queries is unknown, whereas PTQ methods such as OBC (Frantar et al., 2022b) optimize for specific input embeddings chosen from a set of calibration data.
In the rest of this section, we present an adaptation of AQ in the context of neural network compression, by taking into account the distribution over input activations. Specifically, for a linear layer with $d_{in}$ input and $d_{out}$ output features given its weights $\mathbf{W}\in\mathbb{R}^{d_{out}\times d_{in}}$ and a set of calibration inputs $\mathbf{X}\in\mathbb{R}^{d_{in}\times n}$, one seeks for a configuration of quantized weights $\hat{W}$ that optimizes squared error between the output of the original and compressed layer:
@@
In the following, we will assume that $\widehat{\mathbf{W}}$ is quantized using AQ, and adopt standard notation (Martinez et al., 2016). The AQ scheme splits weight rows into groups of $g$ consecutive elements, and represents each group of weights as a sum of $M$ vectors chosen from multiple learned codebooks $C_{1},...C_{M}$, each codebook containing $2^{B}$ vectors (for B-bit codes). A weight is encoded by choosing a single code from each codebook and summing them up. For simplicity, we denote this choice as a one-hot vector $b_{m}$, which results in the following representation for a group: $\sum_{m=1}^{M}C_{m}b_{ijm}$. This is similar to PTQ algorithms such as GPTQ (Frantar et al., 2022a), except for having more complex coding for each group. In contrast, GPTQ can be seen as quantizing each group to a separate code.
To represent the full weight, we simply concatenate:
$\widehat{\mathbf{W}}_{i}{=}\sum_{m=1}^{M}C_{m}b_{i,1,m}\oplus...\oplus\sum_{m=1}^{M}C_{m}b_{i,d_{in}/g,m}$ (3)
Here, $\oplus$ denotes concatenation and $b_{ijm}\in\mathbb{R}^{2^{B}}$ represents a one-hot code for the $i$-th output unit, $j$-th group of input dimensions and $m$-th codebook.
@@
Our algorithm will learn codebooks $C_{m}\in\mathbb{R}^{g\times 2^{B}}$ and the discrete codes represented by one-hot $b\in\mathbb{R}^{d_{out}\times d_{in}/g\times M\times 2^{B}}$. The resulting scheme encodes each group of $g$ weights using $M\cdot B$ bits and further requires $g\cdot 2^{B}\cdot 16$ bits for FP16 codebooks. The compression error to be minimized becomes:
@@
To learn this weight representation, we initialize codebooks $C$ and codes $b$ by running residual K-means following (Chen et al., 2010). Then, we alternate between updating codes $b_{i,j,m}$ and codebooks $C_{m}$ until the loss function (4) stops improving up to the specified tolerance. Since codes are discrete and codebooks are continuous, optimizing them requires two distinct optimization strategies. We describe each of these two optimization phases below.
@@ 3.2 Phase 1: Beam search for codes
During this phase, AQ updates the codes $b_{i,j,m}$ to minimize the MSE objective (4). Similarly to Babenko & Lempitsky (2014); Martinez et al. (2016, 2018), we reformulate the objective in terms of a fully-connected discrete Markov Random Field (MRF) to take advantage of MRF solvers.
To simplify the derivation, let us first consider a special case where there is a single output unit ($d_{out}{=}1)$ and a single quantization group (i.e. $g{=}d_{in}$), to get rid of the concatenation operator: $||\mathbf{W}\mathbf{X}-\sum^{M}_{m=1}C_{m}b_{m}\mathbf{X}||_{2}^{2}$. We rewrite this objective by expanding the squared difference [^2] Best viewed as a tensor generalization of the following equation: $(a-b)^{2}=a^{2}+b^{2}-2ab$:
@@
Here, the $\langle\cdot,\cdot\rangle_{F}$ denotes a Frobenius inner product of two matrices. Intuitively, this is defined as an inner product between the vectors obtained by flattening the two matrices, or equivalently, $\langle\mathbf{A},\mathbf{B}\rangle_{F}=\langle\text{vec}(\mathbf{A}),\text{vec}(\mathbf{B})\rangle=\text{tr}(\mathbf{A}^{T}\mathbf{B})$.
Let us consider the three components of (5) in isolation. First, note that $||\mathbf{W}\mathbf{X}||_{2}^{2}$ is constant in $b$ and can be disregarded during optimization. The second component can be expanded further into pairwise dot products:
@@
In this form, it is clear that one can pre-compute $\mathbf{X}\mathbf{X}^{T}\in\mathbb{R}^{d_{in}\times d_{in}}$. Abusing notation, we denote this type of product as $\left\langle\mathbf{A},\mathbf{B}\right\rangle_{\mathbf{X}\mathbf{X}^{T}}\stackrel{{\scriptstyle\mathsf{def}}}{{=}}\left\langle\mathbf{A}\mathbf{X}\mathbf{X}^{T},\mathbf{B}\right\rangle_{F}$ in future derivations.
Combining all these derivations, equation (5) becomes:
@@
Finally, we generalize this equation to multiple output units ($d_{out}>1$) and quantization groups ($g{\neq}d_{in}$). For $d_{out}>1$, note that the original objective (4) is additive with respect to output units: thus, we can apply (8) independently to each output dimension and add up results. To support multiple input groups ($g{\neq}d_{in}$), we can treat each group as a separate codebook where only the codes for the active group are nonzero. Thus, we need to repeat each codebook $d_{in}/g$ times and pad it with zeros according to the active group.
It is now evident that minimizing (5) is equivalent to MAP inference in a Markov Random Field with $\left\langle\mathbf{W},C_{m}b_{m}\right\rangle_{\mathbf{XX}^{T}}$ as unary potentials and $\left\langle C_{i}b_{i},C_{j}b_{j}\right\rangle_{\mathbf{XX}^{T}}$ as pairwise potentials. While finding the exact optimum is infeasible, prior work has shown that this type of MRF can be solved approximately with beam search or ICM (Besag, 1986).
@@
Since the loss function is additive, changing one code only affects a small subset of loss components. Thus, we can compute the loss function efficiently by starting with a previous loss function (before code replacement), then adding and subtracting the components that changed during this iteration. These few loss components can be computed efficiently by multiplying with $\mathbf{X}\mathbf{X}^{T}$ ahead of beam search. The beam search runs over all $d_{out}$ output units in parallel. This is possible because encoding one output unit does not affect the objective (8) of other units. Note that beam search is not the best known solution to this problem. Modern AQ variants for retrieval (Martinez et al., 2016, 2018) use randomized ICM to find solutions faster. In this study, we chose beam search only because it was easier to implement in high-level frameworks like PyTorch/JAX.
@@ 3.3 Phase 2: Codebook update
In the second phase, we find the optimal codebook vectors $C_{1},...,C_{M}$ that minimize the same squared error as beam search. Note that, if we treat the codes $b$ as constants, minimizing (4) becomes a least squares problem for $C_{m}$. The original AQ algorithm solves this problem in a closed form, relying on the fact that each vector dimension can be optimized independently.
Our problem is more complicated because of the presence of $\mathbf{X}\mathbf{X}^{T}$: the optimal value of one codebook coordinate depends on the values of all other coordinates. In principle, we could still optimize $C_{m}$ in a closed form, but it would require inverting an impractically large matrix. We can also use iterative least squares solvers (e.g. conjugate gradients) specialized for this type of problem.
@@
where $\widehat{\mathbf{W}})$ is the quantized weight matrix from 3, and the $\mathbf{X}\mathbf{X}^{T}$ matrix is pre-computed. We optimize this objective by iterating (non-stochastic) full-batch gradient descent.
For each codebook update phase, our implementation runs 100 Adam steps with learning rate 1e-4. However, we found that the final result is not sensitive to either of these parameters: training with smaller number of steps or learning rate achieves the same loss, but takes longer to converge. In future work, these hyperparameters could be eliminated by switching to dedicated least squares solver for codebooks.
Similarly to most other quantization algorithms, we also learn per-unit scales $s\in\mathbb{R}^{h}$ that are initialized as $s_{i}:=||\mathbf{W}_{i}||_{2}$ and updated alongside codebooks with the same Adam optimizer (line 19 in Algorithm 1).
3.4 Fine-tuning for intra-layer cohesion
So far, our algorithm compresses each weight matrix independently of the rest of the model. However, in practice, quantization errors interact differently between matrices. This issue is especially relevant in the case of for extreme (2-bit) compression, where quantization errors are larger.
@@ 3.4 Fine-tuning for intra-layer cohesion
Table 1: Evaluation of quantized Llama 2 models for 2-2.1 bits per parameter. The table reports perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
7B 16.00 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 2.02 6.93 8.84 64.64 73.50 48.23 62.84 33.62 56.57
QuIP 2.020 222.54 145.03 51.78 54.62 27.25 26.01 19.37 35.81
QuIP# 2.02 8.22 11.01 62.43 71.38 42.94 55.56 28.84 52.23
13B 16.00 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 1.970 5.70 7.59 69.46 75.08 53.44 66.96 38.48 60.68
QuIP 2.00 13.48 16.16 52.80 62.02 35.80 45.24 23.46 43.86
QuIP# 2.011 6.06 8.07 63.38 74.76 51.58 64.06 33.96 57.55
70B 16.00 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 2.070 3.94 5.72 75.93 80.43 61.79 77.68 0.4793 68.75
QuIP 2.01 5.90 8.17 67.48 74.76 50.45 62.16 33.96 57.76
QuIP# 2.01 4.16 6.01 74.11 79.76 60.01 76.85 47.61 67.67
Here, we opt for a middle ground between the two approaches by performing optimization at the level of individual transformer blocks, i.e. groups of 4-8 linear layers [^3] The exact number depends on several of factors including the use of gated GLU activations, group query attention and QKA weight merging. that constitute a single multi-head self-attention, followed by a single and MLP layer. Having quantized all linear layers within a single transformer block, we fine-tune its remaining parameters to better approximate the original outputs of that transformer block by backpropagating through the weight representation (3).
More specifically, we use PyTorch autograd engine to differentiate the $||\text{block}(\mathbf{X}_{block})-\mathbf{Y}_{block}||^{2}$ , where $\mathbf{X}_{block}$ are the inputs activations for that transformer block and $\mathbf{Y}_{block}$ are output activations of $\text{block}(\mathbf{X}_{block})$ recorded prior to quantization. We train the codebooks $C_{m}$, scale vectors $s$ and all non-quantized parameters (RMSNorm scales and biases), while keeping the codes $b_{i,j,m}$ frozen.
Similarly to Section 3.3, we train these parameters using Adam optimizer to minimize the mean squared error against the original block outputs (prior to quantization). This phase uses the same calibration data as for the individual layer quantization. The full procedure is summarized in Alg. 1.
Algorithm 1 Additive Quantization for LLM
@@
While fine-tuning transformer blocks is more expensive than individual linear layers, it is still possible to quantize billion-parameter models on a single GPU. Furthermore, since the proposed algorithm only modifies the few remaining trainable parameters, it uses relatively little VRAM for Adam optimizer states. The fine-tuning procedure typically converges after a few epochs, likely because it starts from a good initial guess. In practice, fine-tuning transformer layers takes up a minority (10-30%) of the total calibration time. However, the entire procedure takes longer than single-pass quantization algorithms such as GPTQ.
Table 2: Evaluation of quantized Llama 2 models for 3-3.1 bits per parameter. The table reports perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
@@
7B 16.00 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 3.04 5.46 7.10 68.35 76.61 54.36 67.76 38.65 61.15
GPTQ 3.00 8.28 10.72 60.30 71.65 45.60 56.48 31.23 53.05
SpQR 2.98 6.21 8.23 63.38 74.27 51.32 62.21 34.13 57.06
QuIP 3.00 - - - - - - - -
13B 16.00 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 3.026 4.83 6.37 67.64 77.80 58.22 73.48 43.60 64.15
GPTQ 3.00 5.87 7.90 64.17 75.14 53.70 68.14 39.16 60.06
SpQR 2.98 5.28 7.07 67.96 76.93 55.72 68.06 38.48 61.43
@@
Table 3: Evaluation of quantized Llama 2 models for 4-4.1 bits per parameter. The table reports perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.

### 06 Feb 2024 v2

@@ 3.1 Overview
We start from the observation that additive quantization (AQ) solves a related problem to post-training quantization (PTQ) (Nagel et al., 2020; Frantar et al., 2022b): both settings assume the existence of a set of “input” vectors, i.e. input data for AQ, and the weight matrix rows for PTQ. The goal is to compress these inputs while preserving dot product similarity, against query vectors (for AQ), and against layer input embeddings (for PTQ). The difference between the two is that AQ assumes that the distribution of queries is unknown, whereas PTQ methods, e.g. (Frantar et al., 2022b), show that it is sufficient to optimize for sample input embeddings from a set of calibration data.
At a high level, we start by solving the following problem: for a linear layer with $d_{in}$ input and $d_{out}$ output features given its weights $\mathbf{W}\in\mathbb{R}^{d_{out}\times d_{in}}$ and a set of calibration inputs $\mathbf{X}\in\mathbb{R}^{d_{in}\times n}$, one seeks for a configuration of quantized weights $\hat{W}$ that optimizes squared error between the output of the original and compressed layer:
@@
In the following, we will assume that $\widehat{\mathbf{W}}$ is quantized using AQ, and adopt standard notation (Martinez et al., 2016). AQ splits weight rows into groups of $g$ consecutive elements, and represents each group of weights as a sum of $M$ vectors chosen from multiple learned codebooks $C_{1},...C_{M}$, each containing $2^{B}$ vectors (for B-bit codes). A weight is encoded by choosing a single code from each codebook and summing them up. We denote this choice as a one-hot vector $b_{m}$, which results in the following representation for a group: $\sum_{m=1}^{M}C_{m}b_{ijm}$. This is similar to PTQ algorithms (Frantar et al., 2022a), except for using much more complex coding per group. To represent the full weights, we simply concatenate:
$\widehat{\mathbf{W}}_{i}{=}\sum_{m=1}^{M}C_{m}b_{i,1,m}\oplus...\oplus\sum_{m=1}^{M}C_{m}b_{i,d_{in}/g,m},$ (2)
where $\oplus$ denotes concatenation and $b_{ijm}\in\mathbb{R}^{2^{B}}$ represents a one-hot code for the $i$-th output unit, $j$-th group of input dimensions and $m$-th codebook.
@@
Our algorithm will learn codebooks $C_{m}\in\mathbb{R}^{g\times 2^{B}}$ and the discrete codes represented by one-hot $b\in\mathbb{R}^{d_{out}\times d_{in}/g\times M\times 2^{B}}$. The resulting scheme encodes each group of $g$ weights using $M\cdot B$ bits and further requires $g\cdot 2^{B}\cdot 16$ bits for FP16 codebooks. The error becomes:
@@
To learn this weight representation, we initialize codebooks $C$ and codes $b$ by running residual K-means as in Chen et al. (2010). Then, we alternate between updating codes $b_{i,j,m}$ and codebooks $C_{m}$ until the loss function (3) stops improving up to the specified tolerance. Since codes are discrete and codebooks are continuous, and we are optimizing over multiple interacting layers, our approach has three phases, described in Algorithm 1 and detailed below.
@@ 3.2 Phase 1: Beam search for codes
First, AQLM updates the codes $b_{i,j,m}$ to minimize the MSE objective (3). Similarly to Babenko & Lempitsky (2014); Martinez et al. (2016, 2018), we reformulate the objective in terms of a fully-connected discrete Markov Random Field (MRF) to take advantage of MRF solvers.
To simplify the derivation, let us first consider a special case of a single output unit ($d_{out}{=}1)$ and a single quantization group (i.e. $g{=}d_{in}$), to get rid of the concatenation operator: $||\mathbf{W}\mathbf{X}-\sum^{M}_{m=1}C_{m}b_{m}\mathbf{X}||_{2}^{2}$. We rewrite this objective by expanding the squared difference:
@@
Above, $\langle\cdot,\cdot\rangle_{F}$ denotes a Frobenius inner product of two matrices. Next, let us consider the three components of Eqn. (4) in isolation. First, note that $||\mathbf{W}\mathbf{X}||_{2}^{2}$ is constant in $b$ and can be ignored. The second component can be expanded further into pairwise dot products:
@@
Thus one can pre-compute $\mathbf{X}\mathbf{X}^{T}\in\mathbb{R}^{d_{in}\times d_{in}}$. We will denote this type of product as $\left\langle\mathbf{A},\mathbf{B}\right\rangle_{\mathbf{X}\mathbf{X}^{T}}\stackrel{{\scriptstyle\mathsf{def}}}{{=}}\left\langle\mathbf{A}\mathbf{X}\mathbf{X}^{T},\mathbf{B}\right\rangle_{F}$ in future derivations. Then, Eqn. (4) becomes:
@@
Finally, we generalize this equation to multiple output units ($d_{out}>1$) and quantization groups ($g{\neq}d_{in}$). For $d_{out}>1$, note that the original objective (3) is additive with respect to output units: thus, we can apply (7) independently to each output dimension and sum up results. To support multiple input groups ($g{\neq}d_{in}$), we can treat each group as a separate codebook where only the codes for the active group are nonzero. Thus, we need to repeat each codebook $d_{in}/g$ times and pad it with zeros according to the active group.
It is now evident that minimizing (4) is equivalent to MAP inference in a Markov Random Field with $\left\langle\mathbf{W},C_{m}b_{m}\right\rangle_{\mathbf{XX}^{T}}$ as unary potentials and $\left\langle C_{i}b_{i},C_{j}b_{j}\right\rangle_{\mathbf{XX}^{T}}$ as pairwise potentials. While finding the exact optimum is infeasible, prior work has shown that this type of MRF can be solved approximately via beam search or ICM (Besag, 1986).
@@
Since the loss function is additive, changing one code only affects a small subset of loss components. Thus, we can compute the loss function efficiently by starting with a previous loss function (before code replacement), then adding and subtracting the components that changed during this iteration. These few loss components can be computed efficiently by multiplying with $\mathbf{X}\mathbf{X}^{T}$ ahead of beam search. The beam search runs over all $d_{out}$ output units in parallel. This is possible because encoding one output unit does not affect the objective (7) of other units. Note that beam search is not necessarily the best solution to this problem. AQ variants for retrieval (Martinez et al., 2016, 2018) use randomized ICM to find solutions faster. In this study, we chose beam search because it was easier to implement in ML frameworks like PyTorch/JAX.
@@ 3.3 Phase 2: Codebook update
In the second phase, we find the optimal codebook vectors $C_{1},...,C_{M}$ that minimize the same squared error as the beam search. If we treat the codes $b$ as constants, minimizing (3) becomes a least squares problem for $C_{m}$. The original AQ algorithm solves this problem in closed form, relying on the fact that each vector dimension can be optimized independently. Our problem is complicated due to the presence of $\mathbf{X}\mathbf{X}^{T}$: the optimal value of one codebook coordinate depends on the values of all others. In principle, we could optimize $C_{m}$ in closed form, but it would require inverting a large matrix, or using iterative least squares solvers (e.g. conjugate gradients) specialized to this problem.
@@
where $\widehat{\mathbf{W}}$ is the quantized weight matrix from 2, and the $\mathbf{X}\mathbf{X}^{T}$ matrix is pre-computed. We optimize this objective by iterating (non-stochastic) full-batch gradient descent.
For each update phase, our implementation runs 100 Adam steps with learning rate 1e-4. However, we found that the final result is not sensitive to either of these parameters: training with smaller number of steps or learning rate achieves the same loss, but takes longer to converge. In future work, these hyperparameters could be eliminated by switching to dedicated least squares solver for codebooks. Similarly to other algorithms, we also learn per-unit scales $s\in\mathbb{R}^{h}$ that are initialized as $s_{i}:=||\mathbf{W}_{i}||_{2}$ and updated alongside codebooks via the same optimizer (line 19 in Algorithm 1).
3.4 Phase 3: Fine-tuning for intra-layer cohesion
So far, our algorithm compresses each weight matrix independently of the rest of the model. However, in practice, quantization errors interact differently between matrices. This issue is especially relevant in the case of extreme (2-bit) compression, where quantization errors are larger.
@@ 3.4 Phase 3: Fine-tuning for intra-layer cohesion
Algorithm 1 AQLM: Additive Quantization for LLMs
@@
Here, we opt for a middle ground by performing optimization at the level of individual transformer blocks, i.e. groups of 4-8 linear layers [^2] This number depends on factors including the use of gated GLU activations, group query attention and QKA weight merging. that constitute a single multi-head self-attention, followed by a single and MLP layer. Having quantized all linear layers within a single transformer block, we fine-tune its remaining parameters to better approximate the original outputs of that transformer block by backpropagating through the weight representation (2).
Concretely, we use the PyTorch autograd engine to differentiate the $||\text{block}(\mathbf{X}_{block})-\mathbf{Y}_{block}||^{2}$ , where $\mathbf{X}_{block}$ are the inputs activations for that transformer block and $\mathbf{Y}_{block}$ are output activations of $\text{block}(\mathbf{X}_{block})$ recorded prior to quantization. We train the codebooks $C_{m}$, scale vectors $s$ and all non-quantized parameters (RMSNorm scales and biases), while keeping the codes $b_{i,j,m}$ frozen. Similarly to Section 3.3, we train these parameters using Adam to minimize the MSE against the original block outputs (prior to quantization). This phase uses the same calibration data as for the individual layer quantization. The full procedure is summarized in Alg. 1.
While fine-tuning blocks is more expensive than individual linear layers, it is still possible to quantize billion-parameter models on a single GPU in reasonable time. Also, since the algorithm only modifies a few trainable parameters, it uses little VRAM for optimizer states. This fine-tuning converges after a few iterations, as it starts from a good initial guess. In practice, fine-tuning transformer layers takes a minority (10-30% or less) of the total calibration time.
Table 1: Evaluation of quantized Llama 2 models for 2-2.8 bits per parameter, with an extra section for higher bitwidth. We report perplexity on WikiText2 (Merity et al., 2016) & C4 (Raffel et al., 2020) and accuracy for zero-shot tasks. The Average accuracy is the mean of 5 zero-shot tasks. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.
@@
7B 16 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 2.02 6.64 8.56 64.17 73.56 49.49 61.87 33.28 56.47
QuIP# 2.02 8.22 11.01 62.43 71.38 42.94 55.56 28.84 52.23
AQLM 2.29 6.29 8.11 65.67 74.92 50.88 66.50 34.90 58.57
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 1.97 5.65 7.51 65.43 76.22 53.74 69.78 37.8 60.59
QuIP 2.00 13.48 16.16 52.80 62.02 35.80 45.24 23.46 43.86
QuIP# 2.01 6.06 8.07 63.38 74.76 51.58 64.06 33.96 57.55
AQLM 2.18 5.41 7.20 68.43 76.22 54.68 69.15 39.42 61.58
AQLM 2.53 5.15 6.80 68.11 76.99 56.54 71.38 40.53 62.71
AQLM 2.76 4.94 6.54 68.98 77.58 57.71 72.90 43.60 64.15
70B 16 3.12 4.97 76.95 81.07 63.99 77.74 51.11 70.17
AQLM 2.07 3.94 5.72 75.93 80.43 61.79 77.68 47.93 68.75
QuIP 2.01 5.90 8.17 67.48 74.76 50.45 62.16 33.96 57.76
QuIP# 2.01 4.16 6.01 74.11 79.76 60.01 76.85 47.61 67.67
Table 2: Evaluation of quantized Llama 2 models for 3-3.1 bits per parameter, with the same metrics as in Table 1.
7B 16 5.12 6.63 67.25 78.45 56.69 69.32 40.02 62.35
AQLM 3.04 5.46 7.08 66.93 76.88 54.12 68.06 38.40 60.88
GPTQ 3.00 8.06 10.61 59.19 71.49 45.21 58.46 31.06 53.08
SpQR 2.98 6.20 8.20 63.54 74.81 51.85 67.42 37.71 59.07
13B 16 4.57 6.05 69.61 78.73 59.72 73.27 45.56 65.38
AQLM 3.03 4.82 6.37 68.43 77.26 58.30 70.88 42.58 64.49
GPTQ 3.00 5.85 7.86 63.93 76.50 53.47 65.66 38.48 59.61
SpQR 2.98 5.28 7.06 67.48 77.20 56.34 69.78 39.16 61.99
@@
4 Experiments
We evaluate the AQLM algorithm in typical scenarios for post-training quantization of modern LLMs. Our evaluation is focused on the Llama 2 model family since it is a popular backbone for fine-tuned models or general LLM applications, e.g. (Dettmers et al., 2023a), and we also present results on Mistral-family models (Jiang et al., 2024). In Section 4.1, we evaluate the full AQ procedure for various Llama 2 models and quantization bit-widths; Section 4.2 presents an ablation analysis for individual AQ components and implementation details.
4.1 Compression quality for modern LLMs
We report perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020) validation sets. We also measure zero-shot accuracy on WinoGrande (Sakaguchi et al., 2021), PiQA (Tata & Patel, 2003), HellaSwag (Zellers et al., 2019), ARC-easy and ARC-challenge (Clark et al., 2018) via the LM Eval Harness (Gao et al., 2021). We broadly follow the evaluation setup of GPTQ (Frantar et al., 2022a).
We consider three main targets in terms of compression ranges: 2-2.8 bits, 3-3.1 bits, and 4-4.1 bits per model parameter. In the results below average bits per parameter takes into account only quantized weights, we do not include parameters kept in floating precision similarly to the related work. The details on the model size estimate are provided in Appendix G. We compare AQ against GPTQ for 3&4 bits (Frantar et al., 2022a), SpQR for 3&4 bits (Dettmers et al., 2023b), QuIP in 2,3 & 4 bits (Chee et al., 2023) and QuIP# for 2&4 bits (Tseng et al., 2023). While GPTQ and SpQR technically support 2-bit quantization, they perform poorly in the 2-3 bit range. For QuIP, we omit results for the 7B model, as we could not achieve competitive performance in this one scenario using the available implementations. (Currently, there is no official implementation of the original QuIP (non-#) for the Llama 2 model.) For QuIP#, we focus on 2 and 4 bit because the available implementation does not yet support 3-bit compression. We calibrate each algorithm using the subset of RedPajama dataset (Computer, 2023), with a sequence length of 4096.
The exact bit-widths for each method are dictated by parameters such as the number of codebooks and code width. We report results for the $2{-}2.8$ and $3{-}3.1$ bitwidth ranges in Tables 1 and 2, respectively. Additional results for $4-4.1$ bits are deferred to Appendix E.2.
The results show that AQLM outperforms the previous best PTQ algorithms across all settings, often by wide margins, especially at high compression. This holds both in terms of PPL across standard validation sets (Wiki-Text2 and C4), and accuracy across zero-shot tasks. Specifically, we observe the highest accuracy gains in the “extreme” 2-2.1 bits per parameter range, where the deviation from the uncompressed model becomes large for all methods.
Table 3: Evaluation of quantized Mixtral (Jiang et al., 2024) models for 2 bits. The table reports perplexity on WikiText2 (Merity et al., 2016) and C4 (Raffel et al., 2020), as well as accuracy for zero-shot tasks. The Average accuracy column is the mean of 5 zero-shot task accuracies. Primary metrics are Wiki2 (PPL), C4 (PPL) and Average accuracy.